OpenStudy (anonymous):

5% less than 28% of x^2 is 10.Then what is 15 less than 28% of x^2.

OpenStudy (anonymous):

5% =10 then 15%=30?

OpenStudy (anonymous):

that is not 15%

OpenStudy (anonymous):

sorry, i was just guessing.. 5 into 15 is 3 x 10.. anyway am wrong;.

OpenStudy (anonymous):

5% of what is less than 28% ??

OpenStudy (anonymous):

Dont know? But thats the complete question...

OpenStudy (anonymous):

Can we take simply 5% ??

OpenStudy (anonymous):

May be I am wrong..

OpenStudy (anonymous):

The options are: 1,2,3,4

OpenStudy (anonymous):

@mukushla can u help?

OpenStudy (anonymous):

Let me try then..

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

i got 5 lol

OpenStudy (anonymous):

How? @nickhouraney

OpenStudy (anonymous):

28% of x^2 => 28/100 x X^2 5% less of that will be=> (28/100 x X^2) -5/100 = 10

ganeshie8 (ganeshie8):

me too 5

OpenStudy (anonymous):

pretty sure its wrong haha

OpenStudy (anonymous):

@ganeshie8 how?

OpenStudy (anonymous):

I got -4.5

ganeshie8 (ganeshie8):

28% of x^2 : .28*x^2 5% of above : .05(.28*x^2) combine : .28*x^2 - .05(.28*x^2) = 10 -----------(1) .28*x^2-15 = ? -------------------(2) .28x^2(1-.05) = 10 .28x^2 = 20 plugin this in (2)

OpenStudy (anonymous):

2.8*x^2-1.5

OpenStudy (anonymous):

I also did the same @ganeshie8

ganeshie8 (ganeshie8):

u also got 5, right ? looks like options are wrong !

OpenStudy (anonymous):

ooooh i see how we did it wrong re-read the question

OpenStudy (anonymous):

i used the equation .28x^2=10

OpenStudy (anonymous):

PLZ check your calculation @ganeshie8

OpenStudy (anonymous):

.28x^2(1-.05) = 10 then, .28x^2=10.5

OpenStudy (anonymous):

got it? @ganeshie8

OpenStudy (callisto):

@nickhouraney How did you get that equation?

OpenStudy (anonymous):

i think it should be: .28*x^2 - .05(.28*x^2) = 10 -----------(1) .28*x^2- .15( .28*x^2) = ? -------------------(2) .28x^2(1-.05) = 10 .28x^2(1-.15) = 20 ???

OpenStudy (anonymous):

hold up callisto let me check something

ganeshie8 (ganeshie8):

ack... yup! .05 i treated as .5 lol

OpenStudy (callisto):

5% less than 28% of x^2 is 10. 28% of x^2 => 0.28x^2 5% less than 28% of x^2 => (1-5%) (0.28x^2) = 0.95 (0.28x^2) 5% less than 28% of x^2 is 10. => 0.95 (0.28x^2) = 10 0.28x^2 = 10/0.95 Then what is 15 less than 28% of x^2. 28% of x^2 = 10/0.95 15 less than 28% of x^2 => 10/0.95 - 15 Doesn't look right at all :|

OpenStudy (anonymous):

nope i have no idea

OpenStudy (anonymous):

Answer will be negative by using that..

OpenStudy (callisto):

That's why it doesn't make sense :(

OpenStudy (anonymous):

Ya..... we all got -4.5

OpenStudy (anonymous):

Something tricky in 5% part.. Of what part is 5% we have to be taken..

OpenStudy (anonymous):

(5% less than 28% of x^2) =10

OpenStudy (anonymous):

Yes but of what part 5% we should take??

OpenStudy (anonymous):

Ya...... I agree with @waterineyes

OpenStudy (anonymous):

could it simply be 23% of x^2 = 10?

OpenStudy (anonymous):

I tried that also @nickhouraney

OpenStudy (anonymous):

but gives no answer

OpenStudy (anonymous):

dang lol

OpenStudy (anonymous):

is there some way of setting up a system of equations with this?

OpenStudy (anonymous):

For that we need to understand the problem at first

OpenStudy (anonymous):

Problem solved I think..

OpenStudy (anonymous):

@waterineyes how?

OpenStudy (anonymous):

My screen is showing: Satellite73 is typing a reply.... ha ha ha..

OpenStudy (anonymous):

\(.95\times .28x=10\implies x=37.594\)

OpenStudy (anonymous):

x^2

OpenStudy (anonymous):

I think its x^2

OpenStudy (anonymous):

maybe it is \(\chi\)

OpenStudy (anonymous):

Ha ha ha..

OpenStudy (anonymous):

what did you get @waterineyes

OpenStudy (anonymous):

\(.85\times .28\times 37.594\) might do it

OpenStudy (anonymous):

You took it seriously @nickhouraney

OpenStudy (anonymous):

:-(

OpenStudy (anonymous):

@nickhouraney i think the point is that to say 5% less than some P is A, means \(.95P=A\)

OpenStudy (anonymous):

in other words, do decrease a number by 5% is the same as multiplying the number by \(.95\)

OpenStudy (anonymous):

But that some number P is : \(\frac{28}{100}x^2\) and \(\frac{28}{100}x\)..

OpenStudy (anonymous):

and not..

OpenStudy (anonymous):

in fact we could say "let 28% of \(x^2\) be \( \zeta\)" so we know \(.95\zeta=10\) or \(\zeta =\frac{10}{.95}\)

OpenStudy (anonymous):

Yeah got it..

OpenStudy (anonymous):

then we want 15% less that \(\zeta\) so we want \(.85\times \frac{10}{.95}\) whatever that is

OpenStudy (anonymous):

but i dont believe in you anymore @waterineyes

OpenStudy (anonymous):

15 not 15%..

OpenStudy (anonymous):

are you serious

OpenStudy (anonymous):

Okay @nickhouraney

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=8.5%2F.95

OpenStudy (anonymous):

@nickhouraney it is not 23% of something because percents do not work that way. the operation is multiplication not addition for example, to find 15% of some number you multiply by \(.15\) to increase a number by 15% you multiply by \(1.15\) to decrease a number by 15% you multiply by \(.85\)

OpenStudy (anonymous):

i got you @satellite73 the wording is very confusing on this one (for me at least)

OpenStudy (anonymous):

it is very confusing, yes. english is not math and percent wording is often ambiguous

OpenStudy (anonymous):

But we have to find: \[\zeta - 15\]

OpenStudy (anonymous):

ooh i see, thought it was 15% less than \(\zeta\)

OpenStudy (anonymous):

That will give you the answer negative.. We all have tried it earlier..

OpenStudy (anonymous):

i also assumed it was 15% but @sauravshakya isnt even here to verifu

OpenStudy (anonymous):

But he said question is right and answer choices are 1, 2, 3 and 4..

OpenStudy (anonymous):

and i refuse to think anymore at this point lol

OpenStudy (anonymous):

\(.95\times \zeta=10\iff \zeta=\frac{10}{.95}\) so i guess you want \(\zeta-15=\frac{10}{.95}-15\) a totally different number

OpenStudy (anonymous):

Yes this will be negative first of all..

OpenStudy (anonymous):

ya so you had it right the first time

OpenStudy (anonymous):

choices are 1,2,3 or 4? then i must have really misread the problem

OpenStudy (anonymous):

Yes choices are 1,2 3, and 4..

OpenStudy (anonymous):

And Suarav is an intelligent guy I don't think he would be in a mood here to fool us..

OpenStudy (anonymous):

???

OpenStudy (anonymous):

Nothing.. Just kidding..

OpenStudy (anonymous):

Did you get the solution how to solve it ??

OpenStudy (anonymous):

That was the question from a engineering college.... That was given to me by a friend

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

Then it is big shame on me..

OpenStudy (anonymous):

I just asked because if you got then do tell me how to solve it.. Ha ha ha..

OpenStudy (anonymous):

yes if you do ever figure it out show us

OpenStudy (anonymous):

I still think that we are unable to convert very first line correctly into mathematics..

OpenStudy (anonymous):

@experimentX can u plz see this

OpenStudy (experimentx):

hold on!!